# Collocation methods for nonlinear differential equations on low-rank manifolds

### v2

@@ 5.1 TT-cross time integration
The time-dependent oblique tangent space projector in (57) is defined at each time $t$ by a set of time-dependent multi-indices $\{\mathcal{I}^{\leq k}(t),\mathcal{I}^{>k}(t)\}$. We assume that such multi-indices are nested (30) at each time $t$ so that the oblique projector has the cross interpolation property (34) at each $t$. The evolution equations defining the TT-cross integrator we obtain by directly evaluating (57) at the multi-indices $\{\mathcal{I}^{\leq k}(t),\mathcal{I}^{>k}(t)\}$ and utilizing the cross interpolation property (34)
@@
where we defined the tensor $G_{Y}(t)=G\left(Y(t),t\right)$. Note that (58) is a set of evolution equations for a subset of entries of the approximate solution $Y(t)\in\mathcal{M}_{\bm{r}}$ that define a TT-cross approximation (54) at any given time $t$, hence the name TT-cross integrator. If $G$ arises from the spatial discretization of a PDE (1) involving differential operators then evaluating $G_{Y}\left(\mathcal{I}^{\leq k-1},:,\mathcal{I}^{>k}\right)$ requires entries of $Y$ at indices adjacent to $\{\mathcal{I}^{\leq k},\mathcal{I}^{>k}\}$. We denote by ($\mathcal{I}^{\leq k}_{(a)}$, ($\mathcal{I}^{>k}_{(a)}$ the union of $\mathcal{I}^{\leq k}$, $\mathcal{I}^{>k}$ with the required adjacent indices and compute the right-hand side tensors in (58) by
@@
Computing (59) is efficient for any nonlinear $G$ that we can evaluate entry-wise, regardless of the low-rank structure in $G$. The evolution equations (58) have a clear computational advantage over the evolution equations of orthogonal dynamical low-rank approximation or step-truncation methods [42], which require a low-rank representation of $G$ to be practical.
@@ 5.1.1 Constructing the low-rank solution in TT format
The (time-dependent) tensor cross interpolation (54) formula
@@
allows one to construct the solution to (57) at any time $t$ from the solution to the equations (58). However, using (60) directly can lead to numerical instability as the time-dependent $r_{k}\times r_{k}$ matrices $Y\left(\mathcal{I}^{\leq k},\mathcal{I}^{>k},t\right)$ can be ill-conditioned. Hereafter we describe a more robust method for computing $Y(t)$ by orthogonalization, omitting the dependence of tensors on $t$ to simplify notation. Take QR-decompositions
@@
Computing $Y$ via (63) instead of (60) yields a more stable numerical algorithm as the matrices $\widehat{\bm{Q}}_{k}=\bm{Q}_{k}^{\langle l\rangle}\left(\bm{l}_{\leq k},:\right)$ are related to the orthogonal bases $\bm{U}_{\leq k}$ from which the multi-index sets $\mathcal{I}^{\leq k},\mathcal{I}^{>k}$ were obtained with the TT-cross-DEIM index selection algorithm and therefore are have smaller condition number than $Y\left(\mathcal{I}^{\leq k},\mathcal{I}^{>k}\right)$. The improvement in condition number is indeed verified by our numerical experiments as shown in Figure 5(d).
@@ 5.2 Interpolatory projector-splitting integrator
Next, we directly applying a splitting integrator to the dynamical low-rank evolution equation (57) to obtain a generalization of the orthogonal projector-splitting integrator introduced in [32] for TTs. Just like the orthogonal projector-splitting integrator, the oblique projector-splitting integrator is robust to small singular values and one step of the integrator can be implemented as an efficient sweeping algorithm over the TT-cores of the solution. Using the expression (14) for the oblique tangent space projector the right-hand side of the dynamical low-rank evolution equation (57) can be written as a sum of $12d-$ 1 terms
@@
in consecutive order to obtain the approximate solution $Y(t_{1})=Y_{d}^{+}(t_{1})$ at time $t_{1}$. Note that the projectors $P_{j}^{+},P_{j}^{-}$ depend on solutions to each substep $Y_{j}^{+}(t)$ or $Y_{j}^{-}(t)$ and thus are time-dependent. In the case of orthogonal tangent space projector-splitting it was shown in [32, Theorem 4.1] that $P_{j}^{+},P_{j}^{-}$ can be kept constant during each substep and each of the differential equations (69) can be solved exactly by updating a single TT-core. The following Theorem is an analogous result for the oblique projector-splitting integrator. We suppress the dependence of the multi-indices $\mathcal{I}^{\leq k},\mathcal{I}^{>k}$ on $t$ although it is assumed that such indices are selected at each time $t$ so that the interpolatory projectors (14) (and therefore the oblique tangent space projector (16)) are well-conditioned.
@@
Each split differential equation in (69) is solved exactly with time-independent projectors $P_{j}^{+}$ and $P_{j}^{-}$ at $Y_{j}^{+}(t_{0})$ and $Y_{j}^{-}(t_{0})$, respectively. Moreover, if $Y_{j}^{+}(t_{0})$ has the TT representation
@@
Proof: The proof follows the same approach as the corresponding result for the orthogonal projector-splitting integrator. First recall that we have shown in the proof of Proposition 3.1 that $P_{\leq j-1}P_{>j}$ maps onto a tangent space of $\mathcal{M}_{\bm{r}}$ at each time $t$. This ensures that $Y_{j}^{+}(t)$ belongs to $\mathcal{M}_{\bm{r}}$ for all $t$ and therefore admits a time-dependent orthogonalized rank-$\bm{r}$ TT decomposition of the form
@@
Similar to the TT-cross evolution equations (58), computing the right-hand side of the differential equation (71) requires evaluating $G_{j}^{+}$ at a subset of $r_{j-1}n_{j}r_{j}$ indices and computing the right-hand side of (72) requires evaluating $G_{j}^{-}$ at a subset of $r_{j}^{2}$ indices. These evaluations are efficient for any $G$ that can be evaluated entry-wise and do not require $G$ to have any low-rank structure. The differential equations (71) and (72) also feature inverses of $r_{j}\times r_{j}$ matrices $U_{\leq j}\left(\mathcal{I}^{\leq j},:\right)$ and $V_{>j}\left(:,\mathcal{I}^{>j}\right)$. These matrices define the interpolatory projectors (14) and we select the multi-indices $\mathcal{I}^{\leq j},\mathcal{I}^{>j}$ at each time $t$, e.g., with the TT-cross-DEIM, to ensure that their condition number is small.
@@ 5.2.1 Sweeping algorithm for interpolatory projector-splitting integrator
One complete step of the splitting integrator can be implemented as an efficient sweeping algorithm that updates the TT-cores from left to right and computes the necessary index sets for oblique projection. To describe one step of the sweeping algorithm from time $t_{0}$ to $t_{1}=t_{0}+\Delta t$, suppose we have an orthogonalized TT decomposition of solution at time $t_{0}$ of the form
@@
from $t_{0}$ to $t_{1}$. The solution $Y_{1}^{+}(t_{1})=K_{1}(t_{1})V_{>1}(t_{0})$ is the starting value $Y_{1}^{-}(t_{0})=Y_{1}^{+}(t_{1})$ for the second equation in (69). We then prepare $Y_{1}^{-}(t_{0})$ for the application of Theorem 5.1 by decomposing $K_{1}(t_{1})=U_{1}(t_{1})\bm{R}_{1}(t_{1})$ to obtain $Y_{1}^{-}(t_{0})=U_{1}(t_{1})\bm{R}_{1}(t_{1})V_{>1}(t_{0})$ where $U_{1}(t_{1})$ is left-orthogonal (see Section 2.1) and compute indices $\mathcal{I}^{\leq 1}(t_{1})=\texttt{DEIM}(\bm{U}_{\leq 1}(t_{1}))$. Now we can apply Theorem 5.1 to solve the second differential equation in (69) by integrating
@@
Iterating these steps until we obtain $Y_{d}^{+}(t_{1})=U_{\leq d-1}(t_{1})K_{d}(t_{1})=Y(t_{1})$ completes one step of the first-order splitting integrator. To take another time step the TT representation of $Y(t_{1})$ must be orthogonalized from right to left, i.e., (6) with $k=1$. During this orthogonalization procedure, the indices $\mathcal{I}^{>j}(t_{1})$ can be computed with the right-to-left TT-cross-DEIM sweep as described in Section 4.1. Similar to the orthogonal projector-splitting integrator, obtaining the second-order Strang projector-splitting integrator is straightforward by composing the Lie-Trotter integrator with its adjoint. In this case the forward sweep described above is performed with step-size $\Delta t/2$ and is then followed by a backward sweep also with step-size $\Delta t/2$. The oblique projector-splitting integrator has the same computational complexity as the TT-cross integrator and in fact the two methods are closely related. The TT-cross method integrates TT-cores and then computes an oblique projection. Meanwhile the projector-splitting method computes the oblique projection first and then integrates TT-cores.
@@ 5.3 Rank-adaptive time integration
To increase the $k$th component of the TT solution rank during integration with the TT-cross integrator we modify Algorithm 2 to sample $\hat{r}_{k}>r_{k}$ indices $\bm{l}_{\leq k}$ from the left singular vectors (45) and $\hat{r}_{k}$ indices $\bm{l}_{>k}$ from right singular vectors in (49) by augmenting the DEIM indices with additional indices selected by another sparse index selection algorithm, e.g., GappyPOD+E [39]. From the $\bm{l}_{\leq k},\bm{l}_{>k}$ we construct $\mathcal{I}^{\leq k},\mathcal{I}^{>k}$ in (46),(50) each with $\hat{r}_{k}$ indices. We then integrate the solution $Y(t)$ forward in time on the manifold $\mathcal{M}_{\hat{\bm{r}}}$ using the equations (58). It is well-known that the solution $Y(t)$ with rank $\bm{r}$ belongs to the boundary of the higher rank manifold $\mathcal{M}_{\hat{\bm{r}}}$ where the tangent space is not well-defined [51]. Nevertheless, the evolution equations (58), which define the interpolatory tangent space projection, are well-defined on the boundary of $\mathcal{M}_{\hat{\bm{r}}}$. These equations allow us to integrate $Y(t)$ forward in time on $\mathcal{M}_{\hat{\bm{r}}}$ thereby increasing the solution rank. To increase the $k$th component of the TT solution rank during integration with the projector-splitting integrator we add new (orthogonal) basis vectors to the TT cores with zero singular value and then sample indices from this augmented basis and apply the projector-splitting integrator to the augmented solution. Once again adding new basis vectors with zero singular values places the approximate solution on the boundary of a higher rank manifol $\mathcal{M}_{\hat{\bm{r}}}$. However the projector-splitting integrator is robust to zero singular values and allows us to integrate the solution off of the boundary of the low-rank manifold.
A simple criterion for determining when to increase the $k$th component of the TT-rank vector is based on the singular values $\{\sigma_{k}(\alpha_{k},t)\}_{\alpha_{k}=1}^{r_{k}}$ of the unfolding matrix $\bm{Y}^{\langle k\rangle}$. We select the rank to ensure that the relative size of the smallest singular value
@@ 6 Numerical examples
To illustrate the key differences between the proposed interpolatory projector-splitting integrator the orthogonal projector-splitting integrator [31] we begin with a two-dimensional example.
@@
We consider the Vlasov-Poisson equation
@@
from [24, Example 4.4] with initial condition $f(x,v,0)=\exp(-20(x^{2}+v^{2}))$, electric field $E(x)=0.5\sin(\pi x)$ and $x\in\Omega_{x}=[-1,1],v\in\Omega_{v}=[-1,1]$. Discretizing $\Omega_{x}$ and $\Omega_{v}$ using $n=64$ points and approximating derivatives with a Fourier pseudo-spectral method [26] we obtain a semi-discrete version of the Vlasov-Poisson equation (88) in the form of a differential equation (2) with $d=2$, i.e., a matrix differential equation. To assess the accuracy of the low-rank tensor integrators, we computed a reference solution $\bm{X}(t)$ by integrating the full matrix differential equation to the semi-discrete equation with the four-stage explicit Runge-Kutta (RK4) method using time step-size $\Delta t=10^{-3}$.
@@
We compared the TT-cross integrator presented in Section 5.1 with a step-truncation method using SVD-based truncation (ST-SVD). For both integrators we used Adams-Bashfort 2 with step-size $\Delta t=10^{-3}$. We utilized the rank-adaptive mechanism described in Section 5.3 for TT-cross with parameter $\epsilon_{l}=10^{-7}$. For the ST-SVD solution we set relative truncation tolerance $\delta=10^{-7}$ at each time step allowing the solution rank to adapt in time accordingly. In Figure 4(b) we plot the rank of the TT-cross and ST-SVD solutions versus time. The rank grows rapidly during time integration which allows us to assess the robustness of the rank-adaptive mechanism for the TT-cross integrator. In Figure 4 we plot the relative error of the TT-cross and ST-SVD solutions in the Frobenius norm versus time. We observe that the TT-cross solution is more accurate than the ST-SVD solution due to the TT-cross solution rank being slightly larger than the rank of the ST-SVD solution at each step. The error of the TT-cross solution remains controlled during time integration, demonstrating the effectiveness of the rank-adaptive mechanism.
Next we compared the interpolatory projector-splitting integrator (i-PS) presented in Section 5.2 with the orthogonal projector-splitting integrator (o-PS) introduced in [31]. In this case ($d=2$), the sweeping algorithm (see Section 5.2.1) for one step of the projector-splitting integrator has three substeps commonly referred to as $K$-, $S$-, and $L$-step. Beginning from the rank-$r$ decomposition of the approximate solution $\bm{Y}(t_{0})=\bm{U}(t_{0})\bm{S}(t_{0})\bm{V}^{\text{T}}(t_{0})$ at time $t_{0}$, these three steps yield low-rank factors for the approximate solution $\bm{Y}(t_{1})=\bm{U}(t_{1})\bm{S}(t_{1})\bm{V}^{\text{T}}(t_{1})$ at time $t_{1}=t_{0}+\Delta t$. The substeps for interpolatory projector-splitting integrator are as follows.
  • 1.
    K-step: update $\bm{U}(t_{0})\to\bm{U}(t_{1})$ and $\bm{S}(t_{0})\to\bm{R}(t_{1})$.
    Compute interpolation indices $\mathcal{J}=\texttt{DEIM}(\bm{V}(t_{0}))$. Then integrate the $n_{1}\times r$ differential equation
    $\frac{d\bm{K}(t)}{dt}=\bm{G}_{\bm{K}}\left(:,\mathcal{J},t\right)\left[\bm{V}\left(\mathcal{J},:,t_{0}\right)\right]^{-\text{T}},\qquad\bm{K}(t_{0})=\bm{U}(t_{0})\bm{S}(t_{0}),$ (89)
    where $\bm{G}_{\bm{K}}(t)=G\left(\bm{K}(t)\bm{V}(t_{0})^{\text{T}},t\right)$ from $t_{0}$ to $t_{1}$, and perform a QR-decomposition $\bm{K}(t_{1})=\bm{U}(t_{1})\bm{R}(t_{1})$.
  • 2.
    S-step: update $\bm{R}(t_{1})\to\tilde{\bm{S}}(t_{1})$.
    Compute interpolation indices $\mathcal{I}=\texttt{DEIM}(\bm{U}(t_{1}))$. Then integrate the $r\times r$ matrix differential equation
    $\frac{d\tilde{\bm{S}}(t)}{dt}=-\left[\bm{U}\left(\mathcal{I},:,t_{1}\right)\right]^{-1}\bm{G}_{\bm{S}}\left(\mathcal{I},\mathcal{J},t\right)\left[\bm{V}\left(\mathcal{J},:,t_{0}\right)\right]^{-\text{T}},\qquad\tilde{\bm{S}}(t_{0})=\bm{R}(t_{1}),$ (90)
    where $\bm{G}_{\bm{S}}(t)=G\left(\bm{U}(t_{1})\bm{S}(t)\bm{V}(t_{0})^{\text{T}},t\right)$ from $t_{0}$ to $t_{1}$.
  • 3.
    L-step: update $\bm{V}(t_{0})\to\bm{V}(t_{1})$ and $\tilde{\bm{S}}(t_{1})\to\bm{S}(t_{1})$.
    Integrate the $n_{2}\times r$ matrix differential equation
    $\frac{d\bm{L}(t)}{dt}=\bm{G}_{\bm{L}}\left(\mathcal{I},:,t\right)^{\text{T}}\left[\bm{U}\left(\mathcal{I},:,t_{1}\right)\right]^{-\text{T}},\qquad\bm{L}(t_{0})=\bm{V}(t_{0})\tilde{\bm{S}}(t_{1})^{\text{T}},$ (91)
    where $\bm{G}_{\bm{L}}(t)=G\left(\bm{U}(t_{1})\bm{L}(t)^{\text{T}},t\right)$ from time $t_{0}$ to $t_{1}$, and perform a QR-decomposition $\bm{L}(t_{1})=\bm{V}(t_{1})\bm{S}(t_{1})^{\text{T}}$.
@@
Figure 4: Low-rank approximations to the solution of the two-dimensional Vlasov-Poisson equation (88). (a) Relative error versus time of TT-cross and ST-SVD solutions. The TT-cross solution was computed using rank-adaptive singular value threshold $\epsilon_{l}=10^{-7}$ and the ST-SVD solution was computed using truncation threshold $\delta=10^{-7}$. (b) Rank versus time of the rank-adaptive TT-cross and ST-SVD solutions. (c) Relative error versus time of solutions computed with interpolatory and orthogonal projector-splitting using rank-adaptive singular value threshold $\epsilon_{l}=10^{-7}$. (d) Rank versus time of the rank-adaptive solutions computed with interpolatory and orthogonal projector splitting integrators.
The interpolatory projector-splitting integrator requires evaluating the tensors $\bm{G}_{\bullet}$ at only a subset of $nr$ indices in the $K$- and $L$-step and $r^{2}$ indices in the $S$-step. These evaluations can be performed efficiently for any $G$ that can be evaluated entry-wise. Meanwhile the corresponding steps of the orthogonal projector-splitting integrator (summarized in [6]) involve inner products between the tensors $\bm{G}_{\bullet}$ that only efficient to compute when $G$ has low-rank structure.
For the interpolatory and orthogonal projector-splitting integrators we used step-size $\Delta t=10^{-3}$ and solved the differential equations in the $K$, $S$, and $L$-steps with RK4. We also used the rank-adaptive mechanism described in Section 5.3 with parameter $\epsilon_{l}=10^{-7}$ for both solutions. In Figure 4(b) we plot the solution ranks versus time. Both solutions have the same rank until approximately $t=0.7$ when the i-PS solution rank becomes slightly smaller than the o-PS solution rank. In Figure 4(a) we plot the relative errors in the Frobenius norm versus time. The error of the i-PS and o-PS solutions is similar until around $t=0.4$, at which point the i-PS solution becomes slightly less accurate. This difference in accuracy is due to the i-PS method computing a quasi-optimal projection onto the tangent space, while the o-PS method computes the optimal projection at each time step. In addition, the slight difference in rank of the solutions also contributes to the difference in accuracy.

### v3

@@ 6.1 TT-cross integrator
The time-dependent interpolatory TT tangent space projector (24) in (3) is defined at each time $t$ by a set of time-dependent multi-indices $\{\mathcal{I}^{\leq k}(t),\mathcal{I}^{>k}(t)\}$. Selecting such with the TT-cross-DEIM ensures that they are nested (38) at each time $t$. Hence the tangent space projector has the cross interpolation property (42) at each $t$. Evaluating (3) at the multi-indices $\{\mathcal{I}^{\leq k}(t),\mathcal{I}^{>k}(t)\}$ and utilizing the cross interpolation property yields evolution equations for the entries of $Y(t)$ defining a TT-cross interpolant
@@
where we defined the tensor $G_{Y}(t)=G\left(Y(t),t\right)$. Equation (65) consists of $\sum_{k=1}^{d}r_{k-1}n_{k}r_{k}$ coupled nonlinear differential equations governing the evolution of a subset of entries in the approximate solution $Y(t)\in\mathcal{M}_{\bm{r}}$, which can be integrated using standard methods. If $G$ arises from the spatial discretization of a PDE (1) involving differential operators then evaluating $G_{Y}\left(\mathcal{I}^{\leq k-1},:,\mathcal{I}^{>k}\right)$ requires entries of $Y$ at indices adjacent to $\{\mathcal{I}^{\leq k},\mathcal{I}^{>k}\}$. Letting ($\mathcal{I}^{\leq k}_{(a)}$, ($\mathcal{I}^{>k}_{(a)}$ denote the union of $\mathcal{I}^{\leq k}$, $\mathcal{I}^{>k}$ with the required adjacent indices, the right-hand side tensors in (65) are computed by
@@
The values of $Y$ at adjacent indices can always be obtained by constructing the low-rank solution $Y(t)$ in the TT format using TT cross interpolation as described below. Computing (66) is efficient for any nonlinear $G$ that we can evaluate entry-wise, regardless of the low-rank structure in $G$. This gives the evolution equations (65) a clear computational advantage over the evolution equations of orthogonal dynamical low-rank approximation or step-truncation methods [42], which require a low-rank representation of $G$ to be practical. We also note that the stiffness of the evolution equations (65) is independent of the singular values of the solution tensor $Y(t)$, unlike other dynamical low-rank methods. However, the stability of constructing the solution in TT format using cross interpolation, which is often needed to evaluate the right-hand side of (65), depends on the condition of the interpolatory projectors obtained through the TT-cross-DEIM index selection as shown below.
@@ 6.1.1 Constructing the low-rank solution in TT format
We can access entries of the approximate solution $Y(t)$ at indices other than the interpolation indices by constructing $Y(t)$ using TT-cross interpolation (62).
@@
Such entries are often needed to evaluate the right-hand side of (65) and the TT-representation (67) is also needed to construct indices for interpolatory projection onto the tangent space. Constructing $Y(t)$ using (67) can lead to numerical instability as the time-dependent $r_{k}\times r_{k}$ matrices $Y\left(\mathcal{I}^{\leq k},\mathcal{I}^{>k},t\right)$ can be ill-conditioned. Hereafter we describe a more robust method for computing $Y(t)$ by orthogonalization, omitting the dependence on $t$ to simplify notation. Take QR-decompositions
@@
Computing $Y$ via (70) instead of (67) yields a more stable numerical algorithm as the matrices $\widehat{\bm{Q}}_{k}=\bm{Q}_{k}^{\langle l\rangle}\left(\bm{l}_{\leq k},:\right)$ are related to the orthogonal bases $\bm{U}_{\leq k}$ from which the multi-index sets $\mathcal{I}^{\leq k},\mathcal{I}^{>k}$ were obtained with the TT-cross-DEIM index selection algorithm and therefore are have smaller condition number than $Y\left(\mathcal{I}^{\leq k},\mathcal{I}^{>k}\right)$. The improvement in condition number is verified by our numerical experiments as shown in Figure 5(d).
@@ 6.2 Interpolatory projector-splitting integrator
Next, we propose a second time integration scheme by directly applying a splitting integrator to the dynamical low-rank evolution equation (3). This method is a direct generalization of the orthogonal projector-splitting integrator introduced in [32] for TTs. As we will see, the oblique projector-splitting integrator satisfies the same desirable properties: it is robust to small singular values, it exactly reproduces low-rank solutions, and one step of the integrator can be implemented as an efficient sweeping algorithm. The derivation of the integrator and the proofs of these results follow the same steps as the orthogonal projector-splitting integrator. Inserting the oblique tangent space projector (24) into the dynamical low-rank evolution equation (3) we see that the right-hand side is a sum of $12d-$ 1 terms
@@
in consecutive order to obtain the approximate solution $Y(t_{1})=Y_{d}^{+}(t_{1})$ at time $t_{1}$. Note that the projectors $P_{j}^{+},P_{j}^{-}$ depend on solutions to each substep $Y_{j}^{+}(t)$ or $Y_{j}^{-}(t)$ and thus are time-dependent. In the case of orthogonal tangent space projector-splitting it was shown in [32, Theorem 4.1] that $P_{j}^{+},P_{j}^{-}$ can be kept constant during each substep and each of the differential equations (76) can be solved exactly by updating a single TT-core. We have an analogous result for the oblique projector-splitting integrator. In the following Theorem we suppress the dependence of the multi-indices $\mathcal{I}^{\leq k},\mathcal{I}^{>k}$ on $t$ although it is assumed that such indices are selected at each time $t$ so that the interpolatory tangent space projector (24) is well-defined.
@@
Each split differential equation in (76) is solved exactly using time-independent projectors $P_{j}^{+}$ and $P_{j}^{-}$ at $Y_{j}^{+}(t_{0})$ and $Y_{j}^{-}(t_{0})$, respectively. Moreover, if $Y_{j}^{+}(t_{0})$ has the TT representation
@@
Proof: The proof follows a similar approach to the analogous proof for the orthogonal projector-splitting integrator. First recall that we have shown in the proof of Proposition 4.1 that $P_{\leq j-1}P_{>j}$ maps onto a tangent space of $\mathcal{M}_{\bm{r}}$ at each time $t$. This ensures that $Y_{j}^{+}(t)$ belongs to $\mathcal{M}_{\bm{r}}$ for all $t$ and therefore admits a time-dependent orthogonalized rank-$\bm{r}$ TT decomposition of the form
@@
Similar to the TT-cross evolution equations (65), computing the right-hand side of the differential equation (78) requires evaluating $G_{j}^{+}$ at a subset of $r_{j-1}n_{j}r_{j}$ indices and computing the right-hand side of (79) requires evaluating $G_{j}^{-}$ at a subset of $r_{j}^{2}$ indices. These evaluations are efficient for any $G$ that can be evaluated entry-wise and do not require $G$ to have any low-rank structure. The differential equations (78) and (79) involve inverses of $r_{j}\times r_{j}$ matrices $U_{\leq j}\left(\mathcal{I}^{\leq j},:\right)$ and $V_{>j}\left(:,\mathcal{I}^{>j}\right)$. These matrices define the interpolatory projectors (22) and we select the multi-indices $\mathcal{I}^{\leq j},\mathcal{I}^{>j}$ with the TT-cross-DEIM at each time $t$ to keep their condition number is small during time integration.
@@ 6.2.1 Sweeping algorithm for interpolatory projector-splitting integrator
One complete step of the interpolatory projector-splitting integrator from time $t_{0}$ to $t_{1}=t_{0}+\Delta t$ can be implemented by sweeping through the cores of $Y$ updating individual cores from $t_{0}$ to $t_{1}$. As we update the TT-cores we also update the multi-index sets $\{\mathcal{I}^{\leq j},\mathcal{I}^{>j}\}$ to ensure the interpolatory projectors remain well-defined. We begin with an orthogonal TT representation of solution at time $t_{0}$ of the form
@@
from $t_{0}$ to $t_{1}$. The solution $Y_{1}^{+}(t_{1})=K_{1}(t_{1})V_{>1}(t_{0})$ is the starting value $Y_{1}^{-}(t_{0})=Y_{1}^{+}(t_{1})$ for the second equation in (76). We then prepare $Y_{1}^{-}(t_{0})$ for the application of Theorem 6.1 by decomposing $K_{1}(t_{1})=U_{1}(t_{1})\bm{R}_{1}(t_{1})$ to obtain $Y_{1}^{-}(t_{0})=U_{1}(t_{1})\bm{R}_{1}(t_{1})V_{>1}(t_{0})$ where $U_{1}(t_{1})$ is left-orthogonal and compute indices $\mathcal{I}^{\leq 1}(t_{1})=\texttt{DEIM}(\bm{U}_{\leq 1}(t_{1}))$. Now we can apply Theorem 6.1 to solve the second differential equation in (76) by integrating
@@
Iterating these steps until we obtain $Y_{d}^{+}(t_{1})=U_{\leq d-1}(t_{1})K_{d}(t_{1})=Y(t_{1})$ completes one step of the first-order splitting integrator. To take another time step the TT representation of $Y(t_{1})$ must be orthogonalized from right to left to obtain an orthogonal representation of the solution at time $t$ in the form of (15) with $k=1$. During this orthogonalization procedure, the indices $\mathcal{I}^{>j}(t_{1})$ can be computed with the right-to-left TT-cross-DEIM sweep as described in Section 5.1. Similar to the orthogonal projector-splitting integrator, obtaining the second-order Strang projector-splitting integrator is straightforward by composing the Lie-Trotter integrator with its adjoint. In this case the forward sweep described above is performed with step-size $\Delta t/2$ and is then followed by a backward sweep also with step-size $\Delta t/2$. The oblique projector-splitting integrator has the same computational complexity as the TT-cross integrator. Just as with the corresponding matrix integrators described in Section 2, the difference between these two integrators is the order in which interpolatory projection and time integration are performed.
@@ 6.3 Rank-adaptive time integration
To increase the $k$th component of the TT solution rank during integration with the TT-cross integrator we modify Algorithm 2 to sample $\hat{r}_{k}>r_{k}$ indices $\bm{l}_{\leq k}$ from the left singular vectors (53) and $\hat{r}_{k}$ indices $\bm{l}_{>k}$ from right singular vectors in (57) by augmenting the DEIM indices with additional indices selected by another sparse index selection algorithm, e.g., GappyPOD+E [39]. From the $\bm{l}_{\leq k},\bm{l}_{>k}$ we construct $\mathcal{I}^{\leq k},\mathcal{I}^{>k}$ in (54),(58) each with $\hat{r}_{k}$ indices. We then integrate the solution $Y(t)$ forward in time on the manifold $\mathcal{M}_{\hat{\bm{r}}}$ using the equations (65). It is well-known that the solution $Y(t)$ with rank $\bm{r}$ belongs to the boundary of the higher rank manifold $\mathcal{M}_{\hat{\bm{r}}}$ where the tangent space is not well-defined [51]. Nevertheless, the evolution equations (65), which define the interpolatory tangent space projection, are well-defined on the boundary of $\mathcal{M}_{\hat{\bm{r}}}$. These equations allow us to integrate $Y(t)$ forward in time on $\mathcal{M}_{\hat{\bm{r}}}$ thereby increasing the solution rank. To increase the $k$th component of the TT solution rank during integration with the projector-splitting integrator we add new (orthogonal) basis vectors to the TT cores with zero singular value and then sample indices from this augmented basis and apply the projector-splitting integrator to the augmented solution. Once again adding new basis vectors with zero singular values places the approximate solution on the boundary of a higher rank manifol $\mathcal{M}_{\hat{\bm{r}}}$. The projector-splitting integrator is robust to zero singular values and allows us to integrate the solution off of the boundary of the low-rank manifold. A simple criterion for determining when to increase the $k$th component of the TT-rank vector is based on the singular values $\{\sigma_{k}(\alpha_{k},t)\}_{\alpha_{k}=1}^{r_{k}}$ of the unfolding matrix $\bm{Y}^{\langle k\rangle}$. We select the rank to ensure that the relative size of the smallest singular value
@@ 7 Numerical examples
We compute a reference solution $X(t)$ for each application below by integrating the differential equation (2) with the four-stage explicit Runge-Kutta (RK4) method using time step-size $\Delta t=10^{-3}$.
@@
We begin with a two-dimensional example ($d=2$) demonstrating the proposed methods on low-rank matrix manifolds described in Section 2. We consider the Vlasov-Poisson equation
@@
from [24, Example 4.4] with initial condition $f(x,v,0)=\exp(-20(x^{2}+v^{2}))$, electric field $E(x)=0.5\sin(\pi x)$ and $x\in\Omega_{x}=[-1,1],v\in\Omega_{v}=[-1,1]$. Discretizing $\Omega_{x}$ and $\Omega_{v}$ using $n=64$ points and approximating derivatives with a Fourier pseudo-spectral method [26] we obtain a semi-discrete version of the Vlasov-Poisson equation (95) in the form of a differential equation (2) with $d=2$, i.e., a matrix differential equation.
@@
We compared the TT-cross integrator presented in Section 6.1 with a step-truncation method using SVD-based truncation (ST-SVD). For both integrators we used Adams-Bashforth 2 with step-size $\Delta t=10^{-3}$. We utilized the rank-adaptive mechanism described in Section 6.3 for TT-cross with parameter $\epsilon_{l}=10^{-7}$. For the ST-SVD solution we set relative truncation tolerance $\delta=10^{-7}$ at each time step allowing the solution rank to adapt in time accordingly. In Figure 4(b) we plot the rank of the TT-cross and ST-SVD solutions versus time and the numerical rank of the reference RK4 solution with singular value threshold $\delta=10^{-7}$, i.e., the number of singular values with relative size larger than $\delta$. The rank grows rapidly during time integration which allows us to assess the robustness of the rank-adaptive mechanism for the TT-cross integrator. In Figure 4(a) we plot the relative error of the TT-cross and ST-SVD solutions in the Frobenius norm versus time. We observe that the TT-cross solution is more accurate than the ST-SVD solution due to the TT-cross solution rank being slightly larger than the rank of the ST-SVD solution at each step. The error of the TT-cross solution remains controlled during time integration, demonstrating the effectiveness of the rank-adaptive mechanism.
Next we compared the interpolatory projector-splitting integrator (i-PS) presented in Section 6.2 with the orthogonal projector-splitting integrator (o-PS) introduced in [31].
@@
Figure 4: Low-rank approximations to the solution of the two-dimensional Vlasov-Poisson equation (95). (a) Relative error versus time of TT-cross and ST-SVD solutions. The TT-cross solution was computed using rank-adaptive singular value threshold $\epsilon_{l}=10^{-7}$ and the ST-SVD solution was computed using truncation threshold $\delta=10^{-7}$. (b) Rank versus time of the rank-adaptive TT-cross and ST-SVD solutions and the numerical rank of the reference RK4 solution with singular value threshold $10^{-7}$. (c) Relative error versus time of solutions computed with interpolatory and orthogonal projector-splitting using rank-adaptive singular value threshold $\epsilon_{l}=10^{-7}$. (d) Rank versus time of the rank-adaptive solutions computed with interpolatory and orthogonal projector splitting integrators and the numerical rank of the reference RK4 solution with singular value threshold $10^{-7}$.
For the interpolatory and orthogonal projector-splitting integrators we used step-size $\Delta t=10^{-3}$ and solved the differential equations in the K-, S-, and L-step with RK4. We also used the rank-adaptive mechanism described in Section 6.3 with parameter $\epsilon_{l}=10^{-7}$ for both solutions. In Figure 4(b) we plot the solution ranks versus time and the numerical rank of the reference RK4 solution with singular value threshold $10^{-7}$. Both solutions have the same rank until approximately $t=0.7$ when the i-PS solution rank becomes slightly smaller than the o-PS solution rank. In Figure 4(a) we plot the relative errors in the Frobenius norm versus time. The error of the i-PS and o-PS solutions is similar until around $t=0.4$, at which point the i-PS solution becomes slightly less accurate. This difference in accuracy is due to the i-PS method computing a quasi-optimal projection onto the tangent space, while the o-PS method computes the optimal projection at each time step. In addition, the slight difference in rank of the solutions also contributes to the difference in accuracy.